Find the distance between (5,-3) and (2,1)

Answer by jim_thompson5910(35256)

d = sqrt((x2-x1)^2+(y2-y1)^2)


d = sqrt((2-5)^2+(1-(-3))^2)


d = sqrt((2-5)^2+(1+3)^2)


d = sqrt((-3)^2+(4)^2)


d = sqrt(9+16)


d = sqrt(25)


d = 5


So exact distance between the two points is 5 units.